OpenStudy (anonymous):

Let A and B be 3x3 matrices with det(A)=2 and det(B)=-3. Find det(A^-1 B)

OpenStudy (loser66):

If A is invertible, then det (A^-) = det(A)^- so that det (A^-) = 1/2 and you have det (AB) = det A*det B apply for det (A^-*B) = det(A^- ) *det B

OpenStudy (anonymous):

thank you very much

OpenStudy (loser66):

ok